Simply and express your answer in a form that is free of radicals in the denominator

3/2+√3-4/1-√2


1
Expert's answer
2021-10-05T13:39:58-0400
"\dfrac{3}{2+\sqrt{3}}-\dfrac{4}{1-\sqrt{2}}=\dfrac{2(2-\sqrt{3})}{4-3}-\dfrac{4(1+\sqrt{2})}{1-2}"

"=4-2\sqrt{3}+4-4\sqrt{2}=8-2\sqrt{3}-4\sqrt{2}"